# Principal stratification with $U$-statistics under principal ignorability

### v2

@@ 5.2 The win-ratio estimands with an intermediate variable
Figure 2: Violin plots of estimators for the win-ratio estimands in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). Horizontal dashed lines represent the true values of the PGCEs.
@@ 6 Data example
We illustrate the proposed methods with a data example from the U.S. Jobs Corps study (Schochet et al.,, 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24. The dataset (named ‘JC’ with $9,240$ observations, available in the R package causalweight) was from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either, the treatment group, i.e., receiving an offer to participate in Job Corps immediately, or, the control group, i.e., not participating in Job Corps until three years later. Follow-up surveys were administered to both groups 24, 36, and 48 months after the random assignment to collect data on the outcomes of interest. As in many open-label randomized controlled trials, substantial noncompliance to treatment assignment as an intermediate outcome occurred 1 year after randomization (15.4% [$857/5,577$] of participants assigned to treatment did not adhere to the assignment, and 50.6% [$1,854/3,663$] assigned to control switched to the treatment condition).
@@
The proposed methods were implemented to examine the causal effect of Job Corps intervention, where the weekly earnings one year after the assignment were selected as the final outcome. As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects. In each data analysis below, we adjust for baseline covariates including baseline characteristics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
6.1 The survivor probability index for continuous outcomes
@@
We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$. In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al., (2009). Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps. For methods based on parametric working models, a logistic model is considered for the employment process within each assignment group, and a log-normal model is fitted for the final earning outcome to induce the pairwise mean outcome function. For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the Super Learner with the random forest and generalized linear models libraries.
@@
Table 1: Estimated SPI for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). The corresponding 95% confidence interval is given in the following parenthesis.
Methods Survivor proability index estimates (standard errors)
@@
6.2 Principal win-ratio for ordinal outcomes
@@
We provide additional analysis of the Job Corps data to illustrate the application of the proposed methods to address ordinal outcomes with noncompliance. In this example, we use the actual treatment status as an intermediate variable, similar to Chen and Flores, (2015). Under the monotonicity assumption, three strata potentially exist, including the compliers (stratum $10$), the never-takers (stratum $00$), and the always-takers (stratum $11$). To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu et al., (2018). Under this setting, we estimate the principal win-ratios within each principal strata similar to the definition in (4). For methods based on parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are estimated via logistic regression. For the debiased machine learning approach, we used the multinomial generalized linear models with the elastic net penalty to estimate the pairwise outcome mean, and the treatment propensity and principal scores are estimated using Super Learning; the estimator is fitted under a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
@@
The estimated results are given in Table 2, where $1,000$ bootstrap replicates were obtained for estimating the standard errors. The estimated proportions of always-takers, compliers, and never-takers in all participants are 50.1%, 34.0%, and 15.9%, respectively. The Jobs Corps intervention assignment positively affects participants’ earnings one year after the assignment for compliers and always-takers, and the effect appears to be more pronounced for compliers. For example, using the debiased machine learning estimator, there is 27% increase in the chances of being in a higher earning category for a complier under treatment than a complier under control. The 95% lower confidence limits obtained from all five methods are strictly above the null value of 1. For the never-takers, however, the effect of Jobs Corps appears more neutral, as the principal win-ratio is estimated to be closer to 1 and three out of five methods (M1, triply robust and debiased machine learning) produce 95% confidence intervals that include the null value of 1. Interestingly, in this reanalysis using a win-ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable. In particular, the statistically significant effect among the always-takers is indicative of a positive psychological effect. That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence having a large chance to move into a higher earning category) compared to those who switched from their initial control assignment.
@@
Table 2: Estimated win-ratios within different compliance strata for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). The corresponding 95% confidence interval is given in the following parenthesis.
@@ 7 Discussion
In this article, we developed principal stratification methods for addressing a general class of PGCE estimands defined with a binary intermediate variable. Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability that are based on the contrast function, we contributed a semiparametric efficient approach for estimating the PGCE estimands by identifying their efficient influence functions and obtaining the ratio $U$-estimators that combine several different nuisance working functions. We showed that the $U$-estimators are triply robust in that they are consistent and asymptotically normal when two of the three nuisance functions are correctly specified. The $U$-estimators become locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric models or by machine learning models with cross-fitting. In light of the increasing interest in pairwise comparison estimands in studies with a composite endpoint (Buyse,, 2010; Bebu and Lachin,, 2015) in biomedical research, our methods can also be useful when there is a need to additionally address a binary intercurrent event, such as treatment noncompliance or death (Lipkovich et al.,, 2022; Mao,, 2024)
@@
The estimands and efficient $U$-estimators were developed to address a binary treatment and a binary intermediate outcome specifically but without restrictions on the final outcome. Although this is a standard setup for principal stratification analysis, additional work is needed to expand the identification conditions and results to address extensions of the PGCE estimands with a continuous treatment and/or a continuous intermediate variable (Antonelli et al.,, 2023). With continuous treatments, the definition of pairwise comparison estimands becomes non-trivial as there are infinitely many contrast functions (based on an infinitesimal increment of the treatment value) that one can specify to define $\tau_{h}$ even in the absence of an intermediate variable. On the other hand, with a continuous intermediate variable, there are infinitely many principal strata that can lead to conceptual and implementation challenges (Schwartz et al.,, 2011). Focusing on a linear contrast function, Lu et al., (2023) have developed nonparametric identification and semiparametric efficient estimators for principal causal effects with a continuous intermediate variable, which may inspire future theoretical development that allows for a more general nonlinear contrast function in the estimands definition.
@@
Finally, as in all principal stratification analyses that achieve point identification, our $U$-estimators for estimating PGCEs depend on the monotonicity and the mean principal ignorability assumptions, that are unverifiable from the observed data alone. In Section S7 of the Online Supplement, we have described a sensitivity function approach to obtain the bias-corrected $U$-estimators for estimating the PGCEs, when either the monotonicity assumption or the mean principal ignorability assumption is violated. The bias-corrected $U$-estimators remain triply robust to potential working model misspecifications and provide means to generate a range of point estimates under assumed departures from the key identification assumptions.

### v3

@@ 5.2 The win ratio estimands with an intermediate variable
Figure 2: Violin plots of estimators for the win ratio estimands in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). Horizontal dashed lines represent the true values of the PGCEs.
@@ 6 Data example
We illustrate the proposed methods with a data example from the U.S. Jobs Corps study (Schochet et al., 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24. The dataset (named ‘JC’ with $9,240$ observations, available in the R package causalweight) was from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, i.e., receiving an offer to participate in Job Corps immediately, or, the control group, i.e., not participating in Job Corps until three years later. Follow-up surveys were administered to both groups 24, 36, and 48 months after the random assignment to collect data on the outcomes of interest. As in many open-label randomized controlled trials, substantial noncompliance to treatment assignment as an intermediate outcome occurred 1 year after randomization (15.4% [$857/5,577$] of participants assigned to treatment did not adhere to the assignment, and 50.6% [$1,854/3,663$] assigned to control switched to the treatment condition).
@@
The proposed methods were implemented to examine the causal effect of Job Corps intervention, where the weekly earnings one year after the assignment were selected as the final outcome. As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects. In each data analysis below, we adjust for baseline covariates, including baseline characteristics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
6.1 The SPI for continuous outcomes
@@
We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$. In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al. (2009). Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps. For methods based on parametric working models, a logistic model is considered for the employment process within each assignment group, and a log-normal model is fitted for the final earning outcome to induce the pairwise mean outcome function. For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear models libraries.
@@
Table 1: Estimated SPI for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). The corresponding 95% confidence interval (CI) is given in the following parenthesis.
Methods SPI estimates
@@
6.2 Principal win ratio for ordinal outcomes
@@
We provide additional analysis of the Job Corps data to illustrate the application of the proposed methods to address ordinal outcomes with noncompliance. In this example, we use the actual treatment status as an intermediate variable, similar to Chen and Flores (2015). Under the monotonicity assumption, three strata potentially exist, including the compliers (stratum $10$), the never-takers (stratum $00$), and the always-takers (stratum $11$). To address the possibility of unemployment status at the time of earning assessment, we create an ordinal outcome with three ordered categories: 0 (unemployment), 0 to 170 per week (hourly wage 4.25 dollars on a 40-hour-per-week basis), and above 170 dollars per week, following the categorization considered by Lu et al. (2018). Under this setting, we estimate the principal win ratios within each principal strata similar to the definition in (4). For methods based on parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are estimated via logistic regression. For the debiased machine learning approach, we used the multinomial generalized linear models with the elastic net penalty to estimate the pairwise outcome mean, and the treatment propensity and principal scores are estimated using Super Learning; the estimator is fitted under a five-fold cross-fitting ($K=5$ and $L=15$) procedure.
@@
The estimated results are given in Table 2, where $1,000$ bootstrap replicates were obtained for estimating the standard errors. The estimated proportions of always-takers, compliers, and never-takers in all participants are 50.1%, 34.0%, and 15.9%, respectively. The Jobs Corps intervention assignment positively affects participants’ earnings one year after the assignment for compliers and always-takers, and the effect appears to be more pronounced for compliers. For example, using the debiased machine learning estimator, there is a 27% increase in the chances of being in a higher earning category for a complier under treatment than a complier under control. The 95% lower confidence limits obtained from all five methods are strictly above the null value of 1. For the never-takers, however, the effect of Jobs Corps appears more neutral, as the principal win ratio is estimated to be closer to 1, and three out of five methods (M1, triply robust and debiased machine learning) produce 95% confidence intervals that include the null value of 1. Interestingly, in this reanalysis using a win ratio estimand, we find that treatment effect heterogeneity may exist across the three compliance strata, and the exclusion restriction assumption often invoked for compliance analysis may be questionable. In particular, the statistically significant effect among the always-takers is indicative of a positive psychological effect. That is, participants assigned to the Jobs Corp intervention may feel encouraged and naturally be more optimistic about future employment opportunities (hence having a large chance to move into a higher earning category) compared to those who switched from their initial control assignment.
@@
Table 2: Estimated win ratios within different compliance strata for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). The corresponding 95% confidence interval is given in the following parenthesis.
@@ 7 Discussion
In this article, we developed principal stratification methods for addressing a general class of PGCE estimands defined with a binary intermediate variable. Under the assumptions of treatment unconfoundedness, monotonicity, and mean principal ignorability that are based on the contrast function, we contributed a semiparametric efficient approach for estimating the PGCE estimands by identifying their efficient influence functions and obtaining the ratio $U$-estimators that combine several different nuisance working functions. We showed that the $U$-estimators are triply robust in that they are consistent and asymptotically normal when two of the three nuisance functions are correctly specified. The $U$-estimators become locally efficient, with their asymptotic variances achieving the semiparametric efficiency lower bounds when all nuisance functions are correctly specified by parametric models or by machine learning models with cross-fitting. In light of the increasing interest in pairwise comparison estimands in studies with a composite endpoint (Buyse, 2010; Bebu and Lachin, 2015) in biomedical research, our methods can also be useful when there is a need to additionally address a binary intercurrent event, such as treatment noncompliance or death (Bornkamp et al., 2021; Lipkovich et al., 2022; Mao, 2024)
@@
The estimands and efficient $U$-estimators were developed to address a binary treatment and a binary intermediate outcome specifically but without restrictions on the final outcome. Although this is a standard setup for principal stratification analysis, additional work is needed to expand the identification conditions and results to address extensions of the PGCE estimands with a continuous treatment and/or a continuous intermediate variable (Antonelli et al., 2023). With continuous treatments, the definition of pairwise comparison estimands becomes non-trivial as there are infinitely many contrast functions (based on an infinitesimal increment of the treatment value) that one can specify to define $\tau_{h}$ even in the absence of an intermediate variable. On the other hand, with a continuous intermediate variable, there are infinitely many principal strata that can lead to conceptual and implementation challenges (Schwartz et al., 2011). Focusing on a linear contrast function, Lu et al. (2023) have developed nonparametric identification and semiparametric efficient estimators for principal causal effects with a continuous intermediate variable, which may inspire future theoretical development that allows for a more general nonlinear contrast function in the estimands definition.
@@
Finally, as in all principal stratification analyses that achieve point identification, our $U$-estimators for estimating PGCEs depend on the monotonicity and the mean principal ignorability assumptions that are unverifiable from the observed data alone. In Section S7 of the Online Supplement, we have described a sensitivity function approach to obtain the bias-corrected $U$-estimators for estimating the PGCEs when either the monotonicity assumption or the mean principal ignorability assumption is violated. The bias-corrected $U$-estimators remain triply robust to potential working model misspecifications and provide means to generate a range of point estimates under assumed departures from the key identification assumptions.
Acknowledgement
Research in this article was partially supported by the United States National Institutes of Health (NIH), National Heart, Lung, and Blood Institute (NHLBI, grant number R00-HL141678). All statements in this report, including its findings and conclusions, are solely those of the authors and do not necessarily represent the views of the NIH.